Please answer this ASAP, Thanks: Suppose an energy market is a monopoly market. Demand is described by P=70−2Q, which means marginal revenue (MR) is described by MR=70-4Q, and supply (MC) is described by P = 3Q . Which of the following statements are true?

The equilibrium monopoly price and quantity are $50 per unit and 10 units, respectively.

The monopoly price is $8 more than the perfectly competitive market price, all else equal.

The transfer (monopoly rent) received by the monopolist is equal to $75.

All of the statements are true.

Two of the three statements are true.
Answer #1

Ans

Here MR=70-4Q

Mc=3Q for equilbrium Mc=MR

70-4Q=3Q

7Q=70

So Q=70/7=10

Price=70-2Q=70-20=50

In case of perfect competition p=mc

So 70-2Q=3Q

70/5=Q

Q=14

P=70-2(14)=42

So right answer is b